# G2D2: Gradient-guided Discrete Diffusion for image inverse problem solving

### v1

@@ A.2 Conditional Generation Using Discrete Diffusion Models as Priors
Appendix B Proofs
Full Statement of Theorem 3.1
In this section, we provide detailed proofs for the main theoretical results presented in the paper.
Theorem B.1 (Full version of Theorem 3.1).
Let $p_{\bm{\alpha}}$ be a distribution with the parameterization given by the decomposition in (5). Then, for any measurements $\mathbf{y}$, the following inequality holds for the KL divergence between the marginal distributions:
$D_{\mathrm{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right],$ (8)
where $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ is the variational marginal distribution parameterized by $\bm{\alpha}$, $q(\mathbf{z}_{0}|\mathbf{y})$ is the true posterior distribution, $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is the variational conditional distribution as defined in (5), $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ is the true conditional distribution, $p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})$ is the marginal distribution at time step $t$, and $T$ is the total number of time steps in the diffusion process.
@@ Appendix B Proofs
To prove the inequality in Theorem 3.1, we start by noting that the KL divergence between the marginal distributions $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ and $q(\mathbf{z}_{0}|\mathbf{y})$ can be bounded by the KL divergence between the joint distributions $p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})$ and $q(\mathbf{z}_{0:T}|\mathbf{y})$:
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{y})\right)$ $\leq D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\ \|\ q(\mathbf{z}_{0:T}|\mathbf{y})\right).$ (9)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\ \|\ q(\mathbf{z}_{0:T}|\mathbf{y})\right)$ $=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}{q(\mathbf{z}_{0:T}|\mathbf{y})}\right]$
$=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right]$
$=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})}{q(\mathbf{z}_{T}|\mathbf{y})}+\sum_{t=1}^{T}\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right]$
$=D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\ \|\ q(\mathbf{z}_{T}|\mathbf{y})\right)+\sum_{t=1}^{T}\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right].$ (10)
In the context of mask-absorbing state diffusion, the distribution $p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})$ is the same as $q(\mathbf{z}_{T}|\mathbf{y})$ because $\mathbf{z}_{T}$ is fully determined by the diffusion process and is independent of $\bm{\alpha}$. Therefore, the first term is zero:
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\ \|\ q(\mathbf{z}_{T}|\mathbf{y})\right)=0.$ (11)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\ \|\ q(\mathbf{z}_{0:T}|\mathbf{y})\right)=\sum_{t=1}^{T}\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right].$ (12)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\ \|\ q(\mathbf{z}_{0:T}|\mathbf{y})\right)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (13)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)$
$\quad+\mathbb{E}_{\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\right)\right]$
$=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)$
$\quad+\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\right)\right].$ (14)
In this equation, the left-hand side represents the KL divergence between $p_{\bm{\alpha}}$ and $q$ at time $t-1$ conditioned on $\mathbf{z}_{t}$, plus the expected KL divergence between their respective conditional distributions of $\mathbf{z}_{0}$. The right-hand side represents the KL divergence between $\tilde{p}_{\bm{\alpha}}$ and $q$ directly conditioned on $\mathbf{z}_{t}$, plus an expected KL divergence over $\mathbf{z}_{0}$.
The crucial observation here is that the last term on the right-hand side is zero. This is because $p_{\bm{\alpha}}$ and $q$ share the same reverse diffusion process when conditioned on $\mathbf{z}_{0}$ and $\mathbf{z}_{t}$, i.e.,
$p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})$ $=q(\mathbf{z}_{t-1}|\mathbf{z}_{0})$
$=q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y}).$ (15)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\right)=0.$ (16)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)$
$=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)$
$\ -\mathbb{E}_{\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\right)\right].$ (17)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)\leq D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right).$ (18)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\ \|\ q(\mathbf{z}_{0:T}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (19)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\ \|\ q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (20)
@@
Full Statement of Lemma 3.2
Lemma B.2 (Full version of Lemma 3.2).
The KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the true posterior $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ can be decomposed into two terms:
$D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=D_{\mathrm{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t})\right)-\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[\log q(\mathbf{y}|\mathbf{z}_{0})\right],$ (21)
where the first term in the right-hand side represents the KL divergence between the variational distribution and the prior distribution without the measurement condition, and the second term is the expected value of the negative log-likelihood $-\log q(\mathbf{y}|\mathbf{z}_{0})$ under the variational distribution.
@@
We begin by considering the KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the true posterior $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$. Given that $\mathbf{z}_{0}$ is a discrete variable, the KL divergence can be expressed as a sum:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}.$ (22)
By applying Bayes’ theorem to the true posterior $q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, we have:
$q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})=\frac{q(\mathbf{z}_{0}|\mathbf{z}_{t})q(\mathbf{y}|\mathbf{z}_{0})}{q(\mathbf{y}|\mathbf{z}_{t})}.$ (23)
Since $q(\mathbf{y}|\mathbf{z}_{t})$ does not depend on $\mathbf{z}_{0}$, it can be treated as a constant and ignored in the KL divergence calculation. Substituting Eq. (23) into Eq. (22), we obtain:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{0}|\mathbf{z}_{t})q(\mathbf{y}|\mathbf{z}_{0})}.$ (24)
@@
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\left[\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{0}|\mathbf{z}_{t})}-\log q(\mathbf{y}|\mathbf{z}_{0})\right].$ (25)
@@
    The first term represents the KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the prior $q(\mathbf{z}_{0}|\mathbf{z}_{t})$:
    $D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q(\mathbf{z}_{0}|\mathbf{z}_{t})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q(\mathbf{z}_{0}|\mathbf{z}_{t})}.$ (26)

### v2

@@ C.2 Conditional Generation Using Discrete Diffusion Models as Priors
Appendix D Proofs
[…]
@@ Appendix D Proofs
To prove the inequality in Theorem 3.3, we start by noting that the KL divergence between the marginal distributions $p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})$ and $q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})$ can be bounded by the KL divergence between the joint distributions $p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})$ and $q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})$:
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})\right)$ $\leq D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\right).$ (9)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\right)$ $=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}{q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})}\right]$
$=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\prod_{t=1}^{T}q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right]$
$=\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})}+\sum_{t=1}^{T}\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right]$
$=D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\right)+\sum_{t=1}^{T}\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right].$ (10)
In the context of mask-absorbing state diffusion, the distribution $p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})$ is the same as $q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})$ because $\mathbf{z}_{T}$ is fully determined by the diffusion process and is independent of $\bm{\alpha}$. Therefore, the first term is zero:
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{T}|\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{T}|\mathbf{y})\right)=0.$ (11)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\right)=\sum_{t=1}^{T}\mathbb{E}_{p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})}\left[\log\frac{p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\right].$ (12)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\right)=\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (13)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)$
$\quad+\mathbb{E}_{\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\right)\right]$
$=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)$
$\quad+\mathbb{E}_{\mathbf{z}_{0}\sim\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\right)\right].$ (14)
In this equation, the left-hand side represents the KL divergence between $p_{\bm{\alpha}}$ and $q_{\text{star}}$ at time t1t-1 conditioned on $\mathbf{z}_{t}$, plus the expected KL divergence between their respective conditional distributions of $\mathbf{z}_{0}$. The right-hand side represents the KL divergence between $\tilde{p}_{\bm{\alpha}}$ and $q_{\text{star}}$ directly conditioned on $\mathbf{z}_{t}$, plus an expected KL divergence over $\mathbf{z}_{0}$.
The crucial observation here is that the last term on the right-hand side is zero. This is because $p_{\bm{\alpha}}$ and $q_{\text{star}}$ share the same conditional posterior when conditioned on $\mathbf{z}_{0}$ and $\mathbf{z}_{t}$, i.e.,
$p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})$ $=q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0})$
$=q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y}).$ (15)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{0},\mathbf{z}_{t},\mathbf{y})\right)=0.$ (16)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)$
$=D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)$
$\ -\mathbb{E}_{\mathbf{z}_{t-1}\sim p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t-1},\mathbf{z}_{t},\mathbf{y})\right)\right].$ (17)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{y})\right)\leq D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right).$ (18)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0:T}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0:T}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (19)
@@
$D_{\text{KL}}\left(p_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{y})\|q_{\text{star-decomp}}(\mathbf{z}_{0}|\mathbf{y})\right)\leq\sum_{t=1}^{T}\mathbb{E}_{\mathbf{z}_{t}\sim p_{\bm{\alpha}}(\mathbf{z}_{t}|\mathbf{y})}\left[D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)\right].$ (20)
@@
[…]
@@
We begin by considering the KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the true posterior $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$. Given that $\mathbf{z}_{0}$ is a discrete variable, the KL divergence can be expressed as a sum:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}.$ (21)
By applying Bayes’ theorem to the true posterior $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$, we have:
$q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})=\frac{q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})}{q_{\text{star}}(\mathbf{y}|\mathbf{z}_{t})}.$ (22)
Since $q_{\text{star}}(\mathbf{y}|\mathbf{z}_{t})$ does not depend on $\mathbf{z}_{0}$, it can be treated as a constant and ignored in the KL divergence calculation. Substituting Eq. (22) into Eq. (21), we obtain:
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})}.$ (23)
@@
$D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\left[\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})}-\log q_{\text{star}}(\mathbf{y}|\mathbf{z}_{0})\right].$ (24)
@@
    The first term represents the KL divergence between the variational distribution $\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})$ and the prior $q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})$:
    $D_{\text{KL}}\left(\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\|q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})\right)=\sum_{\mathbf{z}_{0}}\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})\log\frac{\tilde{p}_{\bm{\alpha}}(\mathbf{z}_{0}|\mathbf{z}_{t},\mathbf{y})}{q_{\text{star}}(\mathbf{z}_{0}|\mathbf{z}_{t})}.$ (25)